OpenStudy (anonymous):

A set of data with a mean of 45 and a standard deviation of 8.3 is normally distributed. Find the value that is +1 standard deviation away from the mean.

OpenStudy (paki):

we have the formula here... z = (x - mu)/sigma +12 = (x - 45)/8.3 solve for "X"...

OpenStudy (paki):

sorry "z" = 1 .... not 12... 12 is by typing mistake...